jon is 13 years older than tony. the sum of their ages is 27. how old are they?
Answer by richwmiller(17219)

a sneaky way
27-13=14
14/2=7
7+13=20
20+7=27
7 and 20 are the ages
Why does this work?
another way.
j=t+13
j+t=27
j-t=13
add
2j=40
j=20
27-20=7=t